OpenStudy (anonymous):

Nathan had an infection, and his doctor wanted him to take penicillin. Because Nathan’s father and paternal grandfather were allergic to penicillin, Nathan has a 75% chance of having the same allergy. The doctor performed a skin test to see whether Nathan would react to it. The test is 98% accurate. If the test predicts that Nathan is not allergic to penicillin, what is the probability after the test that Nathan is not allergic to the drug?

OpenStudy (anonymous):

Answer Choices are: 0.1135 0.2450 0.9423 0.9932

OpenStudy (anonymous):

@Squirrels

OpenStudy (squirrels):

Ok well... the test is 98% accurate. What is 98% in decimal form?

OpenStudy (anonymous):

.98

OpenStudy (squirrels):

Good! Now do you know what to do? It's pretty simple.

OpenStudy (anonymous):

No what should I do now?

OpenStudy (squirrels):

Ok thats fine. Out of 100% accuracy, there is 98%, correct? Well 98% is .98 in decimal form, and 1.00 is 100%. What is the inaccuracy in decimal form?

OpenStudy (anonymous):

What do you mean inaccuracy?

OpenStudy (squirrels):

Well the test is 98% sure that he is allergic to penicillin. That means the test in 2% inaccurate. (.2) Do you get what i mean?

OpenStudy (anonymous):

yes

OpenStudy (squirrels):

Oops sorry i mean to say 0.02*

OpenStudy (anonymous):

0.2450 So would that be the answer?

OpenStudy (anonymous):

@Squirrels

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ybarrap):

Do you know Bayes Theorem?

OpenStudy (anonymous):

No @ybarrap

OpenStudy (ybarrap):

Let Event A be the event that Nathan is allergic to penicillin Let Event B be the event that the penicillin test comes out positive given that Nathan has the allergy We want the probability that Nathan is not allergic to penicillin and the test comes out positive: P(A)*P(B)=(1-.75)*0.98. Does this make sense?

OpenStudy (anonymous):

Yes it does @ybarrap

OpenStudy (ybarrap):

good luck :)

OpenStudy (anonymous):

@M.D.Saxon

OpenStudy (m.d.saxon):

sorry idk sorry

OpenStudy (ybarrap):

Correction: We want the probability that Nathan is not allergic to penicillin and the test comes out positive: $$\large {\left (1-P(A)\right )\times P(B)=(1-.75)*0.98}$$